# DenoiseRotator: Enhance Pruning Robustness for LLMs via Importance Concentration

### v1

@@ Abstract
Pruning is a widely used technique to compress large language models (LLMs) by removing unimportant weights, but it often suffers from significant performance degradation—especially under semi-structured sparsity constraints. Existing pruning methods primarily focus on estimating the importance of individual weights, which limits their ability to preserve critical capabilities of the model. In this work, we propose a new perspective: rather than merely selecting which weights to prune, we first redistribute parameter importance to make the model inherently more amenable to pruning. By minimizing the information entropy of normalized importance scores, our approach concentrates importance onto a smaller subset of weights, thereby enhancing pruning robustness. We instantiate this idea through DenoiseRotator, which applies learnable orthogonal transformations to the model’s weight matrices. Our method is model-agnostic and can be seamlessly integrated with existing pruning techniques such as Magnitude, SparseGPT, and Wanda. Evaluated on LLaMA3, Qwen2.5, and Mistral models under 50% unstructured and 2:4 semi-structured sparsity, DenoiseRotator consistently improves perplexity and zero-shot accuracy. For instance, on LLaMA3-70B pruned with SparseGPT at 2:4 semi-structured sparsity, DenoiseRotator reduces the perplexity gap to the dense model by 58%, narrowing the degradation from 8.1 to 3.4 points. Codes are available at https://github.com/Axel-gu/DenoiseRotator.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [32, 10, 1, 16] have significantly improved their performance in complex reasoning, multimodal processing, and extended context handling. However, their substantial model sizes and computational demands present practical challenges for deployment and inference efficiency. To mitigate these issues, a variety of model compression techniques have been explored, including quantization [14, 28, 3, 37, 38, 27], knowledge distillation [20], and pruning [30, 18, 13, 36].
Among these, pruning stands out as an effective method for reducing parameter count and computational cost by eliminating less important weights. Traditional pruning methods [18] rely on importance score metrics—such as weight magnitude or output sensitivity—to rank parameters and prune those with the lowest scores. Notably, recent methods like SparseGPT [13] and Wanda [36] estimate the impact of removing individual parameters using Taylor approximations, justifying pruning parameters with minimal estimated impact and approximates the output deviation of the layer by summing the importance scores of the pruned parameters.
@@
To overcome this limitation, we propose a fundamentally different perspective. Rather than merely selecting weights to prune, we aim to reshape the distribution of importance score prior to pruning. Specifically, we instantiate this idea through DenoiseRotator, a framework that applies learnable orthogonal transformations to reparameterize weight matrices, leveraging the computational invariance [2] of Transformer architectures. As illustrated in Figure 2, these transformations are trained to rotate the weight matrices in a way that concentrates the importance scores into a smaller subset of parameters, thereby enhancing pruning robustness.
@@
DenoiseRotator Framework: We introduce DenoiseRotator, a concrete instantiation of our importance concentration idea, implemented via learnable orthogonal transformations that leverage the computational invariance [2] of Transformer [39] architectures.
@@
Extensive Empirical Evaluation: We demonstrate the effectiveness of DenoiseRotator on a range of open-source LLMs, including Mistral (7B) [23], LLaMA3 (8B, 70B) [17], and Qwen2.5 (7B, 14B, 32B, 72B) [40]. Our method consistently and significantly improves pruning performance across unstructured and semi-structured sparsity patterns, reducing perplexity and improving accuracy compared to baseline pruning methods.
@@
Figure 1: Overview of the DenoiseRotator framework. The top row illustrates a Transformer [39] layer architecture—used in mainstream models such as LLaMA [17], Mistral [23], and Qwen [40]—that consists of RMSNorm, attention, and feed-forward blocks. In the middle, learnable orthogonal matrices are inserted to rotate the weight matrices, concentrating parameter importance before pruning. The rotated weights are then merged and pruned in the bottom row. In this illustration, linear layers are represented in the $Y=XW$ format.
@@ 2 Background
Neural network pruning is a widely used model compression technique that reduces model size and computational cost by eliminating parameters deemed less important. Existing approaches can be broadly categorized into two paradigms. The first class relies on heuristic metrics to estimate parameter importance and prunes those ranked lowest [30, 18, 13, 36, 42, 9, 25, 19, 43, 11]. The second class formulates pruning as a continuous optimization problem by relaxing the binary pruning mask into a differentiable form, allowing end-to-end training at the cost of significantly increased computational overhead [29, 12].
Computational invariance refers to the property of transformer architecture proposed in SliceGPT [2] that allows certain orthogonal transformations to be applied to their weight matrices without altering the model’s output. Building upon this principle, QuaRot [3] and SpinQuant [28] apply orthogonal rotations to weight matrices and activations to facilitate low-bit quantization by redistributing outlier values more evenly. RotPruner [7] attempts to leverage this property for pruning and reports promising empirical results; however, it lacks theoretical analysis or formal justification for why rotation improves pruning robustness.
@@ 3.1 Enhance pruning robustness via parameter importance concentration
Pruning aims to reduce model size and computational requirements by removing parameters deemed less important. A critical aspect of effective pruning is accurately estimating the importance of each parameter. As mentioned earlier, recent methods such as SparseGPT [13] and Wanda [36] leverage Taylor series to approximate the change in a linear layer output when a parameter is removed, serving as importance metrics to identify and eliminate the least significant parameters.
@@
This metric is equivalent to applying the Optimal Brain Damage (OBD) [25] approach to the optimization problem of pruned weight $\hat{W}$ to minimize the change in the linear layer’s output after pruning. The OBD importance score is the second-order term of Taylor approximation:
@@
By decreasing the total importance of pruned weights, the deviation in the linear layer’s output is minimized, thus preserving model performance and enhancing robustness against degradation caused by pruning. Furthermore, reducing the magnitude of importance score of pruned weights improves the accuracy of Taylor-series-based importance approximations, leading to more precise importance estimation and compensation strategies, such as those used in SparseGPT [13].
@@ 3.2 Reducing information entropy of normalized importance score
Although the objective in Equation 4 provides a principled formulation for importance concentration, it is generally difficult to optimize in practice. On the left-hand side, the selection of the pruning indices $\mathcal{P}$ is a combinatorial problem and is known to be NP-hard [6]. On the right-hand side, sorting-based strategies—commonly used to identify low-importance parameters—are non-differentiable and thus incompatible with gradient-based optimization.
@@ 3.3.1 Integration of orthogonal matrices
This integration leverages structural properties of the Transformer architecture—specifically, the presence of residual connections, layer normalization (which could be transformed to RMS normalization [2]), and attention mechanisms—and is suitable for any model that exhibits similar characteristics.
@@ 3.3.3 Optimization of the orthogonal matrix
where $\mathcal{L}_{i}$ denotes the set of all linear layers within the $i$-th Transformer layer, and $\mathcal{G}$ indexes the normalization groups within each linear layer. Note that we do not apply any additional weighting to the entropy of each group. This is because the number of normalization groups per linear layer is inherently tied to the weight’s shape (e.g., rows or columns), and the value of discrete entropy naturally scales with the number of categories.
@@ 4 Experiment
Models and Tasks We evaluate our method on several recent open-source large language models, including Mistral 7B [23], LLaMA 3 (8B and 70B) [17], and Qwen2.5 (7B, 14B, 32B, and 72B) [40]. Our evaluation encompasses both language generation, measured by perplexity on the WikiText-2 dataset [31], and five widely-used zero-shot tasks: PIQA [5], WinoGrande [35], HellaSwag [41], ARC-e, and ARC-c [8]. Following established practices, we utilize the LM Evaluation Harness [15] with default settings for all evaluations.
Baselines We integrate DenoiseRotator with three pruning methods: the classic Magnitude pruning and two advanced techniques, Wanda and SparseGPT. Both Wanda and SparseGPT require calibration data to estimate input statistics. For consistency, we use the same calibration dataset as in prior work, consisting of 128 sequences with a context length of 2048 tokens, sampled from the C4 training set [34]. We apply a uniform sparsity level across all decoder layers and evaluate two types of sparsity: unstructured 50% sparsity and semi-structured 2:4 sparsity.
@@ 4.1 Main result
For detailed performance results (including the LLaMA 1 and 2 families), please refer to Appendix D.
@@ 4.2 Effectiveness of Entropy Reduction on Pruning Robustness
As shown in Table 3, reducing the entropy of the normalized importance distribution significantly improves model’s performance. At step 100, where average entropy per layer is reduced by 13%, we observe a substantial boost in zero-shot accuracy from 66.88% to 70.54%, along with a decrease in perplexity from 9.567 to 7.701. Further optimization continues to reduce entropy and improve perplexity, although the zero-shot score slightly declines, suggesting overfit in generation task.
@@ 4.3 Inference speedup and overhead analysis
[…]
@@ Appendix B Proof of invariance of total importance
As established in Equation 3, the total importance of parameters can be approximated by the Frobenius norm of the layer output, corresponding to the scenario where all weights are pruned (set to zero). Specifically, according to optimal brain damage [25], we have:
@@ Appendix C QR decomposition reparameterization
Standard gradient descent algorithms and their variants encounter significant challenges when dealing with the orthogonality constraint of matrices. In conventional settings, these algorithms directly update matrix parameters without ensuring that the orthogonality condition is preserved post-update. This can lead to a loss of orthogonality, compromising model integrity and stability. To address this, previous research [28, 21, 7] has frequently adopted Stiefel manifold optimization techniques [22, 26, 4, 24]. These methods are specifically tailored to handle the complex geometric constraints associated with orthogonal matrices, necessitating specialized optimizers and additional computational resources.
@@
Because of these properties, QR decomposition can seamlessly integrate into the computation graph used by automatic differentiation tools. PyTorch [33] provides automatic differentiation support for QR factorization through torch.qr. When QR decomposition is performed, PyTorch automatically manages and propagates gradients with respect to the input matrix $A$. This capability enables gradient-based methods, such as gradient descent, to efficiently update $A$, automatically respecting the orthogonality constraint imposed on $Q$.
@@
Appendix D Performance of pruned model

### v2

@@ Abstract
Pruning is a widely used technique to compress large language models (LLMs) by removing unimportant weights, but it often suffers from significant performance degradation—especially under semi-structured sparsity constraints. Existing pruning methods primarily focus on estimating the importance of individual weights, which limits their ability to preserve critical capabilities of the model. In this work, we propose a new perspective: rather than merely selecting which weights to prune, we first redistribute parameter importance to make the model inherently more amenable to pruning. By minimizing the information entropy of normalized importance scores, our approach concentrates importance onto a smaller subset of weights, thereby enhancing pruning robustness. We instantiate this idea through DenoiseRotator, which applies learnable orthogonal transformations to the model’s weight matrices. Our method can be seamlessly integrated with existing pruning techniques such as Magnitude, SparseGPT, and Wanda. Evaluated on LLaMA3, Qwen2.5, and Mistral models under 50% unstructured and 2:4 semi-structured sparsity, DenoiseRotator consistently improves perplexity and zero-shot accuracy. For instance, on LLaMA3-70B pruned with SparseGPT at 2:4 semi-structured sparsity, DenoiseRotator reduces the perplexity gap to the dense model by 58%, narrowing the degradation from 8.1 to 3.4 points. Codes are available at https://github.com/Axel-gu/DenoiseRotator.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [GPT4, DeepSeek-R1, Claude3.7, gemini2.5] have significantly improved their performance in complex reasoning, multimodal processing, and extended context handling. However, their substantial model sizes and computational demands present practical challenges for deployment and inference efficiency. To mitigate these issues, a variety of model compression techniques have been explored, including quantization [gptq, spinquant, quarot, quipsharp, tseng2024qtip, vptq], knowledge distillation [knowledge_distillation], and pruning [llmpruner, han2016deepcompressioncompressingdeep, sparsegpt, wanda, one-shot].
Among these, pruning stands out as an effective method for reducing parameter count and computational cost by eliminating less important weights. Traditional pruning methods [han2016deepcompressioncompressingdeep] rely on importance score metrics—such as weight magnitude or output sensitivity—to rank parameters and prune those with the lowest scores. Notably, recent methods like SparseGPT [sparsegpt] and Wanda [wanda] estimate the impact of removing individual parameters using Taylor approximations, justifying pruning parameters with minimal estimated impact and approximates the output deviation of the layer by summing the importance scores of the pruned parameters.
@@
To overcome this limitation, we propose a fundamentally different perspective. Rather than merely selecting weights to prune, we aim to reshape the distribution of importance score prior to pruning. Specifically, we instantiate this idea through DenoiseRotator, a framework that applies learnable orthogonal transformations to reparameterize weight matrices, leveraging the computational invariance [slicegpt] of Transformer architectures. As illustrated in Figure 2, these transformations are trained to rotate the weight matrices in a way that concentrates the importance scores into a smaller subset of parameters, thereby enhancing pruning robustness.
@@
DenoiseRotator Framework: We introduce DenoiseRotator, a concrete instantiation of our importance concentration idea, implemented via learnable orthogonal transformations that leverage the computational invariance [slicegpt] of Transformer [transformer] architectures.
@@
Extensive Empirical Evaluation: We demonstrate the effectiveness of DenoiseRotator on a range of open-source LLMs, including Mistral (7B) [mistral7b], LLaMA3 (8B, 70B) [llama3], and Qwen2.5 (7B, 14B, 32B, 72B) [qwen2.5]. Our method consistently and significantly improves pruning performance across unstructured and semi-structured sparsity patterns, reducing perplexity and improving accuracy compared to baseline pruning methods.
@@
Figure 1: Overview of the DenoiseRotator framework. The top row illustrates a Transformer [transformer] layer architecture—used in mainstream models such as LLaMA [llama3], Mistral [mistral7b], and Qwen [qwen2.5]—that consists of RMSNorm, attention, and feed-forward blocks. In the middle, learnable orthogonal matrices are inserted to rotate the weight matrices, concentrating parameter importance before pruning. The rotated weights are then merged and pruned in the bottom row. In this illustration, linear layers are represented in the $Y=XW$ format.
@@ 2 Background
Neural network pruning is a widely used model compression technique that reduces model size and computational cost by eliminating parameters deemed less important. Existing approaches can be broadly categorized into two paradigms. The first class relies on heuristic metrics to estimate parameter importance and prunes those ranked lowest [llmpruner, han2016deepcompressioncompressingdeep, sparsegpt, wanda, loraprune, gblm, obd, obs, ria, prunerzero]. The second class formulates pruning as a continuous optimization problem by relaxing the binary pruning mask into a differentiable form, allowing end-to-end training at the cost of significantly increased computational overhead [l0_regu, maskllm].
Computational invariance refers to the property of transformer architecture proposed in SliceGPT [slicegpt] that allows certain orthogonal transformations to be applied to their weight matrices without altering the model’s output. Building upon this principle, QuaRot [quarot] and SpinQuant [spinquant] apply orthogonal rotations to weight matrices and activations to facilitate low-bit quantization by redistributing outlier values more evenly. RotPruner [rotpruner] attempts to leverage this property for pruning and reports promising empirical results; however, it lacks theoretical analysis or formal justification for why rotation improves pruning robustness.
@@ 3.1 Enhance pruning robustness via parameter importance concentration
Pruning aims to reduce model size and computational requirements by removing parameters deemed less important. A critical aspect of effective pruning is accurately estimating the importance of each parameter. As mentioned earlier, recent methods such as SparseGPT [sparsegpt] and Wanda [wanda] leverage Taylor series to approximate the change in a linear layer output when a parameter is removed, serving as importance metrics to identify and eliminate the least significant parameters.
@@
This metric is equivalent to applying the Optimal Brain Damage (OBD) [obd] approach to the optimization problem of pruned weight $\hat{W}$ to minimize the change in the linear layer’s output after pruning. The OBD importance score is the second-order term of Taylor approximation:
@@
By decreasing the total importance of pruned weights, the deviation in the linear layer’s output is minimized, thus preserving model performance and enhancing robustness against degradation caused by pruning. Furthermore, reducing the magnitude of importance score of pruned weights improves the accuracy of Taylor-series-based importance approximations, leading to more precise importance estimation and compensation strategies, such as those used in SparseGPT [sparsegpt].
@@ 3.2 Reducing information entropy of normalized importance score
Although the objective in Equation 4 provides a principled formulation for importance concentration, it is generally difficult to optimize in practice. On the left-hand side, the selection of the pruning indices $\mathcal{P}$ is a combinatorial problem and is known to be NP-hard [blumensath2008iterative]. On the right-hand side, sorting-based strategies—commonly used to identify low-importance parameters—are non-differentiable and thus incompatible with gradient-based optimization.
@@ 3.3.1 Integration of orthogonal matrices
This integration leverages structural properties of the Transformer architecture—specifically, the presence of residual connections, layer normalization (which could be transformed to RMS normalization [slicegpt]), and attention mechanisms—and is suitable for any model that exhibits similar characteristics.
@@ 3.3.3 Optimization of the orthogonal matrix
where $\mathcal{L}_{I}nvi$ denotes the set of all linear layers within the $i$-th Transformer layer, and $\mathcal{G}$ indexes the normalization groups within each linear layer. Note that we do not apply any additional weighting to the entropy of each group. This is because the number of normalization groups per linear layer is inherently tied to the weight’s shape (e.g., rows or columns), and the value of discrete entropy naturally scales with the number of categories.
@@ 4 Experiment
Models and Tasks We evaluate our method on several recent open-source large language models, including Mistral 7B [mistral7b], LLaMA 3 (8B and 70B) [llama3], and Qwen2.5 (7B, 14B, 32B, and 72B) [qwen2.5]. Our evaluation encompasses both language generation, measured by perplexity on the WikiText-2 dataset [wikitext2], and five widely-used zero-shot tasks: PIQA [piqa], WinoGrande [winogrande], HellaSwag [hellaswag], ARC-e, and ARC-c [allenai:arc]. Following established practices, we utilize the LM Evaluation Harness [lm-eval-harness] with default settings for all evaluations.
Baselines We integrate DenoiseRotator with three pruning methods: the classic Magnitude pruning and two advanced techniques, Wanda and SparseGPT. Both Wanda and SparseGPT require calibration data to estimate input statistics. For consistency, we use the same calibration dataset as in prior work, consisting of 128 sequences with a context length of 2048 tokens, sampled from the C4 training set [c4]. We apply a uniform sparsity level across all decoder layers and evaluate two types of sparsity: unstructured 50% sparsity and semi-structured 2:4 sparsity.
@@ 4.1 Main result
For detailed performance results (including the LLaMA 1 and 2 families), please refer to Appendix G.
@@ 4.2 Effectiveness of Entropy Reduction on Pruning Robustness
As shown in Table 3, reducing the entropy of the normalized importance distribution significantly improves model’s performance. At step 100, where average entropy per layer is reduced by 13%, we observe a substantial boost in zero-shot accuracy from 66.88% to 70.54%, along with a decrease in perplexity from 9.567 to 7.701. Further optimization continues to reduce entropy and improve perplexity. A comprehensive hyperparameter analysis, detailing optimization steps and learning rates, is available in Appendix D.
@@ 4.3 Inference speedup and overhead analysis
DenoiseRotator introduces a pair of orthogonal matrices at the beginning and end of each layer. Since the orthogonal matrices of adjacent layers can be merged, the actual overhead is limited to storing an additional shape matrix $(\text{hidden\_size},\text{hidden\_size})$ per layer, along with a matrix multiplication during inference. For example, in LLaMA-3-8B, this results in approximately 0.5 billion additional parameters (i.e., $4096\times 4096\times 32$). This overhead is relatively small, accounting for only about 6.7% of the original model size. To further address overhead concerns, we explore the potential of block diagonal orthogonal matrices as a reduction strategy in Appendix E.
@@ 6 Limitations
7 Acknowledgements
This work was supported by the China STI 2030-Major Projects under Grant No. 2021ZD0201500, and it was also supported in part by the Beijing Nova Program.
@@ Appendix B Proof of invariance of total importance
As established in Equation 3, the total importance of parameters can be approximated by the Frobenius norm of the layer output, corresponding to the scenario where all weights are pruned (set to zero). Specifically, according to optimal brain damage [obd], we have:
@@ Appendix C QR decomposition reparameterization
Standard gradient descent algorithms and their variants encounter significant challenges when dealing with the orthogonality constraint of matrices. In conventional settings, these algorithms directly update matrix parameters without ensuring that the orthogonality condition is preserved post-update. This can lead to a loss of orthogonality, compromising model integrity and stability. To address this, previous research [spinquant, ostquant, rotpruner] has frequently adopted Stiefel manifold optimization techniques [james1976topology, Li2020Efficient, becigneul2018riemannian, geoopt2020kochurov]. These methods are specifically tailored to handle the complex geometric constraints associated with orthogonal matrices, necessitating specialized optimizers and additional computational resources.
@@
Because of these properties, QR decomposition can seamlessly integrate into the computation graph used by automatic differentiation tools. PyTorch [pytorch] provides automatic differentiation support for QR factorization through torch.qr. When QR decomposition is performed, PyTorch automatically manages and propagates gradients with respect to the input matrix $A$. This capability enables gradient-based methods, such as gradient descent, to efficiently update $A$, automatically respecting the orthogonality constraint imposed on $Q$.
@@
Appendix D Hyperparameter Sensitivity Analysis
A thorough analysis of hyperparameter sensitivity is essential for understanding how different hyperparameter choices impact the final performance of a model. In the main paper, we relied on a fixed set of hyperparameters, such as using 2000 steps in the rotation training process. To enhance the rigor and reproducibility of our research, we conducted comprehensive experiments on the LLaMA-3-8B model, exploring the effects of various hyperparameter configurations on the final results.
In our experiments, each row presents the perplexity and average zero-shot accuracy corresponding to different learning rates and training steps. The setup involves utilizing SparseGPT with 50% unstructured pruning, using a calibration dataset consisting of 128 samples of length 2048 from the C4 dataset. All experiments were performed on an A100 80G GPU.
Table 8: Results across various learning rates and training steps
Steps 0.1 0.01 0.001 0.0001
Perplexity Zero-shot Perplexity Zero-shot Perplexity Zero-shot Perplexity Zero-shot
100 7.75 69.52 7.70 70.54 7.96 69.95 8.52 67.83
200 7.70 70.31 7.64 68.93 7.82 70.04 8.30 68.59
400 7.63 69.84 7.62 70.12 7.66 70.31 8.16 69.37
800 7.61 70.25 7.61 70.21 7.72 69.55 7.94 70.77
2000 7.64 69.82 7.60 69.58 7.67 70.36 7.85 70.11
4000 7.60 70.46 7.60 69.59 7.63 70.37 7.78 69.59
These results indicate: 1. When the learning rate is less than or equal to 0.001, the efficiency is low due to the small learning rate. 2. With a learning rate of 0.01, training converges around 2000 steps, and excessive steps increase training costs. 3. Compared to the baseline without the importance concentration mechanism, all configurations show significant improvements as training progresses. 4. Zero-shot task accuracy fluctuates around 70%, with a variation of approximately 1%, not showing a clear linear relationship with perplexity.
Appendix E Trade-off with Block Diagonal Orthogonal Matrices
Block diagonal orthogonal matrices provide an effective approach to reduce both training and inference overhead of DenoiseRotator. By leveraging the structural efficiency of block diagonal matrices, this method achieves a balanced optimization of performance and resource utilization.
In our experiments with the LLaMA-3-8B model, we evaluated these benefits using SparseGPT with 50% unstructured pruning over 2000 steps and a learning rate of 0.01. Dense orthogonal matrices were replaced with block diagonal matrices, consisting of block_num orthogonal submatrices arranged along the diagonal. Each submatrix has dimensions of $\frac{\text{hidden\_size}}{\text{block\_num}}$. This approach reduces both spatial and computational complexity to $\frac{1}{\text{block\_num}}$ of the original requirements.
Table 9: Performance and Computational Costs with Different Block Configurations
Block Number Perplexity Zero-Shot Accuracy Time Cost per Step (s) Entropy
1 7.597 69.58 0.124 384128
2 8.024 68.68 0.088 410816
4 8.544 68.47 0.076 428160
8 8.882 67.51 0.072 440512
These results highlight the trade-offs between performance and computational demands across different block configurations, illustrating the capacity of block diagonal orthogonal matrices to optimize these aspects effectively.
Looking ahead, potential enhancements could involve: 1. Varying the dimensions of each block within the block diagonal matrix (as opposed to maintaining uniform block sizes in this section’s experiments) to further improve performance. 2. Incorporating permutation matrices, which have low computational complexity, to combine with block diagonal orthogonal matrices, thus emulating dense orthogonal matrices. This approach might involve arranging highly correlated dimensions or parameters adjacently for enhanced efficacy.
Appendix F Compatibility with LoRA for Fine-Tuning Pruned Models
In this section, we explore the use of LoRA [hu2022lora] to fine-tune pruned models. By freezing rotation matrices during fine-tuning, we maintain consistent gradient flow, aligning with scenarios where rotations are not applied, thus ensuring stability throughout the training process.
DenoiseRotator Configuration
  • Steps: 100
  • Learning Rate: 0.01
  • Model: LLaMA-3-8B
  • Pruning Method: Wanda, 2:4 semi-structured
Fine-tuning Configuration
  • Method: LoRA
  • Alpha: 32.0
  • Dropout: 0.1
  • LoRA-r: 8
  • Dataset: 4096 samples of length 2048 from the WikiText2 train set
  • Learning Rate: 2e-4
  • Weight Decay: 1e-2
  • Optimizer: Adam
  • Learning Rate Scheduler: Linear
  • Warm-up Steps: 400
  • Note: Rotation matrices are kept frozen during fine-tuning
Table 10: Performance of Fine-Tuning with LoRA on Pruned Models
Method Perplexity Zero-Shot Accuracy
Dense 6.14 72.72
Wanda 25.19 51.03
+ DenoiseRotator 14.31 60.67
+ Finetune 9.32 61.08
These results illustrate that additional fine-tuning with LoRA further reduces perplexity and improves accuracy. This suggests that fine-tuning post-denoising is an effective strategy for boosting performance, while confirming that the rotation matrices introduced do not destabilize the training process.
Appendix G Performance of pruned model